# Koopman Spectral Analysis from Noisy Measurements based on Bayesian Learning and Kalman Smoothing

### v1

@@ II-A Koopman operator theory
where ${\bm{x}}\in{\mathcal{M}}\subseteq\mathbb{R}^{n}$ represents the state vector and ${\mathcal{M}}$ represents the state space. The flow induced by this system is denoted as $S^{t},t>0$, i.e., ${\bm{x}}(t)=S^{t}({\bm{x}}(0))$. The Koopman operator $U^{t}:\mathcal{F}\to\mathcal{F}$ is a linear operator acting on the observable functions of the states, i.e., $g\in\mathcal{F}:\mathcal{M}\to\mathbb{C}$, which is defined as
$U^{t}g=g\circ S^{t}.$ (2)
@@ III-B Matrix identification and sample estimation
  • The $i$-th E-step: We employee Kalman smoothing to estimate $\bm{x}_{k}$, where $p(\bm{x}_{k}|Y;\bm{\theta}^{i})\sim\mathcal{N}(\bm{m}_{k}^{i},P_{k}^{i})$, where $\bm{\theta}^{i}$ denotes the $i$th estimation of $\bm{\theta}$. Initializing $\bm{\mu}_{1}^{i}$ and $\Sigma_{1}^{i}$, we have
@@ IV Numerical experiments
In this section, we conduct numerical experiments to demonstrate the robustness of the KBK method proposed in Section III in the presence of noise. The primary objective of these experiments is to evaluate the method’s efficacy in two critical aspects. First, the KBK method applies on nonlinear systems to illustrate the accuracy of the approximated Koopman eigenvalues under varying noise conditions. Second, the system states are estimated using the KBK method, assessing the KBK method’s ability to recover the system’s dynamic behavior in noisy environments.
To highlight the superior performance of the KBK method in noisy environments, we performed a comparative analysis with traditional methods, including DMD method, total least squares DMD (TDMD) [14], forward-backward DMD (fbDMD) [13], and Kalman filter DMD (KF-DMD) [15]. To ensure a fair comparison, same data were applied across all methods in each experiment. Moreover, we utilize the relative error of approximated eigenvalue, which is defined as follows
$E_{1}=\frac{\|\bm{\lambda}_{\text{approx}}-\bm{\lambda}_{\text{true}}\|_{2}}{\|\bm{\lambda}_{\text{true}}\|_{2}},$ (30)
where $\bm{\lambda}_{\text{approx}}$ represents the vector of approximated Koopman eigenvalues, and $\bm{\lambda}_{\text{true}}$ represents the vector of associated theoretical Koopman eigenvalues. Additionally, the error of estimated states is evaluated by root mean squared error (RMSE), i.e.,
$E_{2}=\sqrt{\frac{1}{N}\sum_{k=0}^{N-1}\left(\hat{x}_{1}(kT_{s})-x_{1}(kT_{s})\right)^{2}}$ (31)
where $N$ is the total number of the states, $\{\hat{x}_{1}(kT_{s})\}_{k=0}^{N-1}$ are estimated values and $\{x_{1}(kT_{s})\}_{k=0}^{N-1}$ are associated truth.
We consider three types of nonlinear systems for evaluation: a system with complex eigenvalues, a system with purely imaginary eigenvalues, and a system with real Koopman eigenvalues.
@@ IV-A Nonlinear system with real Koopman eigenvalues
This dynamical system can be represented by a finite-dimensional linear model using the observable functions ${g_{1}({\bm{x}})=x_{1},g_{2}({\bm{x}})=x_{2},g_{3}({\bm{x}})=x_{1}^{2}}$. Under this representation, the Koopman eigenvalues are $-1$ and $-2$, both of which lie on the real axis. Due to all eigenvalues being negative real numbers, the system’s state variables converge monotonically and exponentially to the equilibrium point $(0,0)$, lacking oscillatory and periodic characteristics.
In this experiment, we utilize $30$ snapshot pairs with additive Gaussian noise, where the noise variance $\sigma^{2}$ is set to $\{10^{-4},10^{-3},10^{-2},10^{-1}\}$. In the KBK method, we select the observable space $\mathrm{span}\{g({\bm{x}}),\ldots,U^{3T_{s}}g({\bm{x}})\}$, where $T_{s}=0.2$s. The corresponding approximation results are depicted in Fig. 1, demonstrating the impact of varying noise levels on the accuracy of the methods’ performance. Fig. 1a illustrates the complex plane of the eigenvalue approximation results using the five methods under a noise variance of $10^{-2}$. The theoretical Koopman eigenvalues of the nonlinear system is indicated by black square markers, while scatter points in various colors represent the approximations computed by different methods. It shows that the imaginary part of the approximated result for all methods are close to zero, aligning well with the true values. However, in terms of the real part, the eigenvalues estimated by the KBK method are significantly closer to the true values, as indicated by the black squares, compared to the other methods. This highlights the superior accuracy of the KBK method in capturing the growth and decay rates of the system’s trajectory, making it more reliable for analyzing the dynamical behavior.
@@ IV-C Nonlinear system with complex eigenvalues
Fig. 3c illustrates the state estimation error associated with various methods under different noise levels, with the error $E_{2}$ calculated using (31). The results indicate that when the noise variance is below $10^{-1}$, the KBK method significantly outperforms the other methods in terms of estimation accuracy. This superiority can be attributed to the Kalman smoothing integrated into the KBK method, which enhances the filtering of sampled data, thereby improving estimation precision. Conversely, although the KFDMD method also employs Kalman filtering, it demonstrates poor performance in data estimation due to inaccuracies in the model, which is directly derived from noisy data. Moreover, the efficacy of the KFDMD method is highly dependent on prior knowledge of the noise characteristics. Nevertheless, as the noise level increases, the estimation error of the KBK method similarly escalates. When the noise variance reaches $10^{-1}$, the reconstruction error of the KBK method becomes comparable to that of TDMD; however, it continues to outperform the other three methods.
@@ V Conclusion
This study proposes a robust method for approximating the Koopman spectrum from noisy measurements. By integrating Bayesian learning with Kalman smoothing, it constructs a statistical framework that effectively mitigates the impact of measurement noise within the Koopman operator framework. The method iteratively enhances performance by updating both state estimation and operator identification, thereby improving overall accuracy. A notable strength of this approach lies in its robustness across varying noise level for Koopman spectrum approximation and states reconstruction, even though the noise level is unknown. This resilience under varying noise conditions is thoroughly validated through extensive comparative experiments.
However, there are important challenges remaining for future investigation. Notably, for certain dynamical systems, the effectiveness of most Koopman spectral analysis methods is strongly dependent on the choice of the initial point, which renders these methods particularly sensitive to noise in the data. This sensitivity can lead to significant variations in spectral approximations. In future work, we plan to focus on developing more robust spectral analysis methods to reduce the dependence on initial conditions, making them more reliable and effective for a broader range of dynamical systems.

### v2

@@ II-A Koopman operator theory
where ${\bm{x}}\in{\mathcal{M}}\subseteq\mathbb{R}^{n}$ represents the state vector and ${\mathcal{M}}$ represents the state space. The flow induced by this system is denoted as $S^{t},t>0$, i.e., ${\bm{x}}(t)=S^{t}({\bm{x}}(0))$. The Koopman operator $U^{t}:\mathcal{F}\to\mathcal{F}$ is a linear operator acting on the observable functions of the states, i.e., $g\in\mathcal{F}:\mathcal{M}\to\mathbb{C}$, which is defined as:
$U^{t}g=g\circ S^{t},$ (2)
where $\circ$ denotes the composition of two functions.
@@ II-B Problem formulation
In this paper, we propose a new approach to extract the Koopman spectral properties of (1) and estimate the states $\{{\bm{x}}(kT_{s})\}_{k=0}^{N-1}$ from its noisy measurements simultaneously.
@@ III-B Matrix identification and sample estimation
  • The $i$-th E-step: We employee Kalman smoothing to estimate $\bm{x}_{k}$, where $p(\bm{x}_{k}|Y;\bm{\theta}^{i})\sim\mathcal{N}(\bm{m}_{k}^{i},P_{k}^{i})$, where $\bm{\theta}^{i}$ denotes the $i$th estimation of $\bm{\theta}$. Initializing the mean and covariance of $\bm{x}_{0}$ as $\bm{\mu}_{1}^{i}$ and $\Sigma_{1}^{i}$, respectively, we have
@@ IV Numerical experiments
In this section, we conduct comprehensive numerical experiments to validate the noise robustness of the proposed Koopman Bilinearization with Kalman Filtering (KBK) method introduced in Section III. These experiments are designed to rigorously assess the method’s performance in two primary areas: (1) the accurate estimation of Koopman eigenvalues across varying noise levels, and (2) the effective reconstruction of system states under noisy conditions. Additionally, we briefly discuss the experience of selection for the number of data points $N$ and the dimension of the basis $M$ in experiments as a guideline.
To establish a meaningful benchmark, the KBK method is compared with four established approaches: standard Dynamic Mode Decomposition (DMD), total least squares DMD (TDMD) [14], forward-backward DMD (fbDMD) [13], and Kalman filter-enhanced DMD (KF-DMD) [16]. All methods are evaluated on the same datasets with controlled noise levels to ensure a fair comparison.
For quantitative evaluation, two error metrics are employed. First, the relative eigenvalue error is defined as
$E_{1}=\frac{\|\bm{\lambda}_{\mathrm{approx}}-\bm{\lambda}_{\mathrm{true}}\|_{2}}{\|\bm{\lambda}_{\mathrm{true}}\|_{2}},$ (30)
where $\bm{\lambda}_{\mathrm{approx}}\in\mathbb{C}^{n}$ are the estimated Koopman eigenvalues and $\bm{\lambda}_{\mathrm{true}}\in\mathbb{C}^{n}$ represent the theoretically derived eigenvalues. Second, the accuracy of state estimation is measured via the root mean squared error (RMSE):
$E_{2}=\sqrt{\frac{1}{N}\sum_{k=0}^{N-1}\left(\hat{x}_{1}(kT_{s})-x_{1}(kT_{s})\right)^{2}},$ (31)
where $N$ denotes the total number of sampled states, and $\{\hat{x}_{1}(kT_{s})\}_{k=0}^{N-1}$ and $\{x_{1}(kT_{s})\}_{k=0}^{N-1}$ are the estimated and true states, respectively.
To provide a comprehensive assessment, three representative nonlinear systems are examined: one with real eigenvalues, one with purely imaginary eigenvalues, and one with complex eigenvalues. These systems are chosen to highlight the versatility and noise robustness of the KBK method in handling a diverse range of dynamical behaviors.
@@ IV-A Nonlinear system with real Koopman eigenvalues
This dynamical system can be represented by a finite-dimensional linear model using observable functions ${g_{1}({\bm{x}})=x_{1},g_{2}({\bm{x}})=x_{2},g_{3}({\bm{x}})=x_{1}^{2}}$. Under this representation, the Koopman eigenvalues are $-1$ and $-2$, both of which lie on the real axis. Due to all eigenvalues being negative real numbers, the system’s state variables converge monotonically and exponentially to the equilibrium point $(0,0)$, lacking oscillatory and periodic characteristics.
In this experiment, we utilize $30$ snapshot pairs with additive Gaussian noise, where the noise variance $\sigma^{2}$ is set to $\{10^{-4},10^{-3},10^{-2},10^{-1}\}$. In the KBK method, we select the observable space $\mathrm{span}\{g({\bm{x}}),\ldots,U^{3T_{s}}g({\bm{x}})\}$, where $T_{s}=0.2$s. The corresponding approximation results are depicted in Fig. 1, demonstrating the impact of varying noise levels on the accuracy of the methods’ performance.
Fig. 1a illustrates the complex plane of the eigenvalue approximation results using the five methods under a noise variance of $10^{-2}$. The theoretical Koopman eigenvalues of the nonlinear system is indicated by black square markers, while scatter points in various colors represent the approximations computed by different methods. It shows that the imaginary part of the approximated result for all methods are close to zero, aligning well with the true values. However, in terms of the real part, the eigenvalues estimated by the KBK method are significantly closer to the true values, as indicated by the black squares, compared to the other methods. This highlights the superior accuracy of the KBK method in capturing the growth and decay rates of the system’s trajectory, making it more reliable for analyzing the dynamical behavior.
@@ IV-C Nonlinear system with complex eigenvalues
Fig. 3c illustrates the state estimation error associated with various methods under different noise levels, with the error $E_{2}$ calculated using (31). The results indicate that when the noise variance is below $10^{-1}$, the KBK method significantly outperforms the other methods in terms of estimation accuracy. The results show that when the noise variance is below $10^{-1}$, the KBK method significantly outperforms others in estimation accuracy, thanks to the Kalman smoothing that improves data filtering and precision. In contrast, the KFDMD method, despite using Kalman filtering, performs poorly due to model inaccuracies stemming from noisy data, and its effectiveness heavily relies on prior knowledge of noise characteristics. As noise levels rise, the KBK method’s estimation error increases, and at a variance of $10^{-1}$, its reconstruction error becomes comparable to TDMD, though it still outperforms the other methods.
@@
The KBK method, which iteratively updates between Kalman smoothing and system identification, incurs higher computational costs than one-shot methods like DMD or TDMD. However, for the systems considered in this section, the overhead remains acceptable, with the entire procedure typically completing in about 5 seconds. This modest increase in computation time is well justified by the substantially improved accuracy of the estimated eigenvalues. For parameters, increasing $N$ improves the accuracy but eventually reaches a plateau. Raising $M$ generally reduces the MSE, though it may introduce minor oscillations in the error curve. Both adjustments increase computational cost and the number of iterations. A practical approach is to start with smaller values and gradually increment them until the additional gains become negligible.
@@ V Conclusion
This study proposes a robust method for approximating the Koopman spectrum from noisy measurements. By integrating Bayesian learning with Kalman smoothing, it constructs a statistical framework that effectively mitigates the impact of measurement noise within the Koopman operator framework. The method enhances performance by iteratively updating both state estimation and operator identification. A notable strength of this approach lies in its robustness across varying noise level for Koopman spectrum approximation and states reconstruction, even though the noise level is unknown. This resilience under varying noise conditions is thoroughly validated through extensive comparative experiments.
However, for certain dynamical systems, the effectiveness of most Koopman spectral analysis methods is strongly dependent on the choice of the initial point. This sensitivity can lead to significant variations in spectral approximations. In future work, we aim to develop more robust methods that reduce this dependence, improving reliability for a wider range of dynamical systems.